# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ Abstract
A general theory of stochastic decision forests reconciling two concepts of information flow – decision trees and refined partitions on the one hand, filtrations from probability theory on the other – is constructed. The traditional “nature” agent is replaced with a one-shot lottery draw that determines a tree of a given decision forest, while each “personal” agent is equipped with an oracle providing updates on the draw’s result and makes partition refining choices adapted to this information. This theory overcomes the incapacity of existing approaches to extensive form theory to capture continuous time stochastic processes like Brownian motion as outcomes of “nature” decision making in particular. Moreover, a class of stochastic decision forests based on paths of action indexed by time is constructed, covering a large fraction of models from the literature and constituting a first step towards an approximation theory for stochastic differential games in extensive form.
@@
[1]organization=Technische Universität Berlin, addressline=Straße des 17. Juni 136, city=10623 Berlin, country=Germany
@@ Introduction
This is the first of a series of papers whose objective is to provide a unified theory of stochastic games and decision problems in extensive form. Representing a decision problem in extensive form means specifying it in terms of what I call “extensive form characteristics”, namely, the flow of information about past choices and exogenous events, as well as the set of local choices available to decision makers. Although the classical theory developed by von Neumann and Morgenstern in [45] and Kuhn in [30, 31] is bound to strong finiteness assumptions, the concept itself is very general and encompasses a large amount of applications: in a series of papers among which are [3, 4, 2] and which the monograph [5] builds upon, Alós-Ferrer and Ritzberger develop an abstract and utterly general theory of extensive form games and decision problems and give a concrete order-theoretic characterisation of its own boundaries. It follows that beyond these boundaries, the notions of strategy, randomisation, outcome, and equilibrium – if used in respect to extensive form characteristics of the decision problem at hand – have no rigorous decision-theoretic meaning.
@@
Hence, any decision problem and game exhibiting extensive form characteristics lies within these boundaries, or is at least some sort of limit of objects within these boundaries. The inclusion, or at least the precise meaning of this limit, permits to rigorously state the decision-theoretic meaning of pure and behaviour strategies, outcomes, equilibrium etc. for the given decision problem. This opens a new perspective on problems in continuous time in particular, a sphere in which the subtlety of this issue has been known for a long time (see, e.g., [42, 44] and the references therein). Based on the maximality result in [44], systematically underlined in [4, 2], that urges to restrict outcomes to certain piece-wise constant paths, in [1] a rigorous extensive form foundation of continuous time games and decision problems with such outcomes is proposed for the first time.
@@
How can a stochastic game exhibiting “extensive form characteristics” be modelled? The standard reply from game and decision theory is: introduce a “nature” agent executing a behaviour strategy (see, e.g., [3, section 2.2.3], or [21, 5], going back to Shapley’s work [39]). But here is a fundamental problem in the continuous time case since many relevant modelling applications, from engineering (see [15, 26] and the references therein) over economics (see, e.g., [35, 24]) and finance (see [17, 34, 9, 32]) to reinforcement learning (see, for example, [25]), demand exogenous noise given by stochastic processes whose paths are all but piece-wise constant, the paradigmatic case being Brownian motion (see, for instance, [34, 29, 14] for an overview over both theory and applications). By no means this assumes the “world to be Brownian” or the like; however, such models are used extensively, be it because it appears plausible to permit unpredictably small time lags until the next exogenous information revelation, be it because certain quantities arguably evolve in decision paths of low regularity, be it for mathematical and computational convenience. This creates the fundamental issue of a lack of a proper decision-theoretic foundation of a large class of relevant models. Nevertheless, one may certainly argue that these models exhibit “extensive form characteristics” without being strictly captured within the Alós-Ferrer–Ritzberger framework ([5]).
@@
Despite this, stochastic control and differential game theory based on stochastic analysis have found a pragmatic way to deal with a continuous time decision problem’s “extensive form characteristics” (see, e.g., [14, 13]). While extensive form theory is based on graph theory and refined partitions, stochastic analysis describes exogenous information by means of filtrations on a given measurable space of scenarios, which are strictly less restrictive objects in that $\sigma$-algebras need not be generated by partitions of the sample space. This works as if a one-shot lottery draw selected a scenario $\omega$ in the beginning without being communicated and then, as time progresses, more and more properties of $\omega$ became known to the decision maker in question. This can be, for instance, the knowledge about the realised path of a Brownian motion. From a decision-theoretic point of view, there is no loss in working with this weaker structure as concerns exogenous information.
@@
As to the actual decision makers, however, the stochastic analysis-inspired approach is not compatible with the extensive form paradigm in a strict sense, because it does not explain strategies, outcomes, (subgame-perfect) equilibria in terms of a decision tree-like object and choices locally available at moves. In this approach, typically, so-called “strategies” are defined as stochastic processes satisfying certain measurability property (e.g., progressive measurability) and such that a given stochastic differential equation depending on them in a non-anticipating way has a unique solution in some sense. This contradicts the problem’s own extensive form characteristics in its strict sense because of the ex post restriction of the strategy space which would imply that the availability of a choice depends on future choices (see the introduction in [1] for a detailed discussion of this point). The “outcome” is taken to be the solution to the differential equation and “equilibria” (or “optimal controls”) are defined with implicit reference to the dynamic programming principle, but without extensive form underpinning. Apart from fundamentally compromising the use of equilibrium concept as such, this lack can moreover cause some confusion, for instance, as to the decision-theoretic meaning of these “equilibria” (for example, concerning the notions of “closed” and “open loop” “equilibria” in stochastic differential games, see the discussion in [13, p. 72–76]) and to the precise definition of subgames (for example, in stochastic timing games, see [35] for a discussion).
@@
Thus, there is scope for an extensive form theory that models exogenous information via filtration-like objects, but endogenous decision making via decision tree-like objects and local choices, therefore synthesising both approaches in a prolific way to what I call stochastic extensive forms. This is the aim of this project whose first part is this paper. This theory can be used to construct extensive form games with general stochastic processes as noise. What is more, it can be used to make precise in what sense stochastic control and differential games can be approximated by stochastic extensive form decision problems, for instance for the stochastic timing game, which is typically not formulated in extensive form (see in particular [35], where instead strategic form games are stacked according to a notion of “consistency” which is justified a priori by an analogy to discrete time and a posteriori, in [43], by a “discrete time with an infinitesimal grid” approximation argument reminiscent of [41, 42]). While a central motivation for this project stems from problems in continuous time, the theory is not restricted to this domain. For example, stochastic extensive form theory is independent of the representation of noise via a “nature” agent’s virtual decision making, and in that sense truly stochastic.
@@
The first step in this endeavour, which is taken in the present paper, is to build a theory of stochastic decision forests. These can be thought of as forests of rooted decision trees such that each tree corresponds to exactly one exogenous scenario, and equipped with a structure of similarity across trees identifying moves in a way that is consistent with the tree structure. Decision trees as graph-theoretical objects are the traditional base model for extensive form games and decision problems, but as pointed out in [3] their refined partitions-based representation in terms of the set of all maximal chains exhibits not only strong duality properties, but makes sequential and dynamic decision making amenable to the traditional decision-theoretic paradigm of choice under uncertainty. In this refined partitions approach, just as Savage’s acts map states to outcomes ([38]), a strategy maps any move to a choice given by a set of outcomes still possible at this move from the agent’s perspective. This is the way uncertainty about future choices (of oneself or opponents), that is, endogenous information is dealt with. The way uncertainty about exogenous noise alias exogenous information is modelled in this paper, however, is crucially different. It is the above-mentioned similarity structure, called random moves, that can be equipped with a filtration-like object dynamically revealing information about the realised exogenous scenario, allowing for general stochastic processes as drivers of noise without running into outcome generation problems along that dimension. The adaptedness of strategies can be based on a concept of adapted choices that is introduced in the final section.
@@
The fundamental motivation for this theory arises in a large class of concrete applications. Thus, its presentation is tightly accompanied by examples. While some of them remain rather pedagogical, a general class of stochastic decision forests based on paths of action indexed by time is also constructed. A small set of readily verifiable axioms allows for many different time regimes, many different specifications of the outcomes (e.g. paths of timing games, also in the case of the scenario-dependent expiration of certain options), and general stochastic noise. Of course, it also includes the deterministic case. To the best of the author’s knowledge, this theoretical unification of action path based decision problems with extensive form characteristics within one framework is a second new contribution of this paper in its own right.
@@
Before ending the introduction, three remarks remain to be made. First, this is the first of three papers, hence, like in a piano sonata in three movements, the first part is self-contained and addresses an independent research problem in its own right, but it clearly prepares and is closely related to the other two papers and they must be thought together. Second, as pointed out by Aumann (see [6]), game theory arguably is “interactive decision theory”, while decision theory is, in a trivial sense, single-player game theory. Still, game theory and decision theory are not the same, the first being more about the “interactive” and the latter more about the “decision”. This paper is more concerned with the latter, hence it primarily employs the corresponding phraseology. But, if the context permits it, “game” will be used instead of “decision problem” or vice versa, and so on, and the reader is presumed to be able to deduce the meaning from the context. Finally, the relevant proofs of all new theorems, propositions, lemmata, and claims in examples, for all sections, can be found in the corresponding subsection of the appendix.
@@
  • 6.
    $|M|$ = if $M$ is not a number, then this is the cardinality of the set $M$;
  • 7.
@@
  • 8.
    $\mathscr{A}|_{D}=\{A\cap D\mid A\in\mathscr{A}\}$, for any $\sigma$-algebra $\mathscr{A}$ and any $D\in\mathscr{A}$;
  • 9.
    $\mathscr{A}_{1}\vee\mathscr{A}_{2}$ = the smallest $\sigma$-algebra on $\Omega$ containing both $\mathscr{A}_{1}$ and $\mathscr{A}_{2}$, given a set $\Omega$ and sets $\mathscr{A}_{1},\mathscr{A}_{2}\subseteq\mathcal{P}(\Omega)$.
@@ 1 Decision forests
The basic object of classical extensive form decision and game theory is the decision tree. In the stochastic generalisation presented here, nature does not act as an agent taking decisions dynamically, but simply includes a device “randomly” selecting the decision tree the “personal” agents follow during their decision making process. This implies that we consider decision forests, rather than decision trees. Although this term is not new (as testified by [36]), we consider it in the context of abstract decision theory, and more precisely within the – decision-theoretically natural – refined partitions framework, and in the aim of making this framework amenable to exogenous noise in the sense of general probability theory. This framework has originally been developed for trees, rather than forests, pioneered in [45, section 8], and developed in much more generality in [3, 4] and subsequent papers, much of which is covered in the monograph [5]. In this section, we fix some order-theoretic language, present a definition and interpretation of decision forests within the refined partitions approach, and show that decision forests are nothing but forests of decision trees.
@@ 1.1 Order- and graph-theoretic conventions
The elements of $\mathcal{F}$ are called connected components of $(F,\geq)$. The maximum of a rooted tree $(T,\geq)$ is called the root. The roots of a forest $(F,\geq)$ are the roots of its connected components. A decision forest (decision tree) is a rooted forest (tree, respectively) $(F,\geq)$ such that all $x,y\in F$ with $x\neq y$ can be separated by some maximal chain $c\subseteq F$, that is, $c\cap\{x,y\}$ is a singleton. A move in a decision forest $(F,\geq)$ is a non-terminal node $x\in F$.
@@ 1.2 Decision forests
The following definition is, formally, a transcription of the characterisation of (a subclass of) “game trees” in [3, theorem 3]. So, although, the object is formally not new, we look at it from a slightly different perspective which is why it is recalled here. First, it emphasises the fact that there can be multiple connected components, but restricts the attention to those [3]-“game trees” whose connected components are rooted. Second, the terminology in this text is different also in that it insists on the purely decision-theoretic aspect. What is called “decision forest (or tree) over a set” and “decision path” here, respectively, is called “game tree“ and “play” in [3]. Third, while [3]-“game trees” are defined via certain set-theoretic properties which can be rather easily verified in applications, and are then characterised via the so-called “representation by plays”, the present text perceives the decision-theoretic essence of decision forests over sets rather as being the duality between outcomes and nodes expressed by that representation, and thus uses it as the definition.
@@
Let $V$ be a set. A decision forest over $V$ is a $V$-poset $F$ such that:
@@
$F$ is called decision tree over $V$ iff, in addition, for all $x,y\in F$ there is $z\in F$ with $z\supseteq x\cup y$.
@@
Following [3], $V$ can be seen as the set of possible outcomes of the game, and the forest $(F,\supseteq)$ specifies how, as the decision problem is tackled dynamically, the set of realisable outcomes becomes smaller and smaller. Thus, elements of $F$ can be interpreted as nodes. See [3] for handy criteria on the set $F$ characterising the situation of it being a decision forest over $V$: essentially, this corresponds to $F$ being a “game tree” over $V$ in the sense of [3, definition 4] such that all connected components of $(F,\supseteq)$ have maximal elements.
Condition (1) clarifies that the predecessors of any node $x\in F$ can be totally ordered and contain a root, and therefore constitute an account of the past with a beginning, a history. The set $W$ of maximal chains can be interpreted as the set of decision paths. Condition (2) says that possible outcomes and decision paths are in one-to-one correspondence, in such a way that a node is contained in some decision path iff the corresponding outcome is contained in that node. Even more is true: according to [3, theorem 3, corollary 2]) and remark 1.2, the bijection $f$ is uniquely determined. More precisely:
@@
Let $F$ be a decision forest over some set $V$ and $f\colon V\to W$ be a map as in definition 1.3, where $W$ is the set of maximal chains in $(F,\supseteq)$. Then,
@@
Let $F$ be a decision forest over a set $V$. Then the poset $(F,\supseteq)$ is a decision forest.
@@
Then $F$ defines a decision forest over $V$, and $(F,\supseteq)$ is order-isomorphic to $(F_{0},\geq)$.
Rephrasing the “representation by plays” from [3], we call $F$ the representation by decision paths of $(F_{0},\geq)$. The definition of a decision forest over a set requires essentially that the operations from the two preceding proposition are, up to isomorphism, inverse to each other. See [3, section 4] for more details on this.
Thus, there is a rigorous sense in that both of the mentioned descriptions are equivalent. While the graph-theoretical approach is graphically more convenient, the refined partitions approach is in line with the classical decision theory of choice under uncertainty, in the spirit of Savage (see the introduction of [3] for a discussion). We note that, although phrased differently and with a different aim, all three preceding propositions and the duality concepts constitute one of the essential innovations of [3].

### v2

@@ Abstract
A general theory of stochastic decision forests is developed to bridge two concepts of information flow: decision trees and refined partitions on the one side, filtrations from probability theory on the other. Instead of the traditional “nature” agent, this framework uses a single lottery draw to select a tree of a given decision forest. Each “personal” agent receives dynamic updates from an own oracle on the lottery outcome and makes partition-refining choices adapted to this information. This theory addresses a key limitation of existing approaches in extensive form theory, which struggle to model continuous-time stochastic processes, such as Brownian motion, as outcomes of “nature” decision making. Additionally, a class of stochastic decision forests based on time-indexed action paths is constructed, encompassing a wide range of models from the literature and laying the groundwork for an approximation theory for stochastic differential games in extensive form.
@@
[1]organization=Institut für Mathematik, Technische Universität Berlin, addressline=Straße des 17. Juni 136, city=10623 Berlin, country=Germany
@@ Introduction
This paper is the first in a series aimed at developing a unified theory of stochastic games and decision problems in extensive form. Representing a decision problem in extensive form means specifying it in terms of what the author suggests calling “extensive form characteristics”, namely, the flow of information about past choices and exogenous events, along with the set of choices available to decision makers. Although classical theory, as established by von Neumann and Morgenstern in [52] and furthered by Kuhn in [34, 35], relies on strong finiteness assumptions, the concept itself is very general and broadly applicable. In a series of papers, including [3, 4, 2], and in the monograph [5], Alós-Ferrer and Ritzberger develop an abstract, highly general theory of extensive form games and decision problems and give a concrete order-theoretic characterisation of its own boundaries. Beyond these boundaries, the notions of strategy, outcome, and equilibrium lose rigorous decision-theoretic meaning when applied to the extensive form characteristics of a given decision problem.
@@
Hence, any decision problem and game exhibiting extensive form characteristics lies within these boundaries or is, at least, some sort of limit of objects within these boundaries. The inclusion, or the precise meaning of this limit, allows us to rigorously determine the decision-theoretic meaning of strategies, outcomes, equilibrium, etc. for the given decision problem. This opens a new perspective on problems in continuous time in particular – a domain where the subtlety of this issue has long been recognised (see, e.g. [49, 51] and the references therein). Based on the maximality result in [51], systematically emphasised in [4, 2], which advocates restricting outcomes to certain piecewise constant paths, [1] introduces, for the first time, a rigorous extensive form foundation of continuous-time games and decision problems involving such outcomes.
@@
Having said that, an important question remains to be addressed: How can the “extensive form characteristics” of stochastic games, decision problems exposed to randomness, or those endowed with randomisation devices be modelled? The standard approach in game and decision theory is to introduce a “nature” agent executing a behaviour strategy (described in, e.g. [3, Subsection 2.2.3] or [21, 5], with roots in the works of Shapley [46] and Harsanyi [27, 28, 29]). However, a fundamental problem arises in the continuous-time case: many relevant modelling applications – from economics (see, e.g. [41, 24]) and finance (see [17, 38, 9, 36]) to engineering (see [15, 30] and the references therein) and reinforcement learning (see, for example, [25]) – require exogenous noise represented by stochastic processes whose paths are anything but piecewise constant, with Brownian motion serving as a paradigmatic case (see, for instance, [38, 33, 14] for an overview of both theory and applications). By no means this implies “the world to be Brownian” or any similar assumption; rather, such models are widely used, be it because it appears reasonable to allow for unpredictably small time lags until the next exogenous information revelation, be it because certain quantities arguably evolve in decision paths of low regularity, be it for mathematical and computational convenience. One can certainly argue that these models exhibit “extensive form characteristics” without fitting into the Alós-Ferrer–Ritzberger framework ([5]). This creates a fundamental issue: the lack of a rigorous decision-theoretic foundation of a large class of relevant models.
@@
Despite this, stochastic control and differential game theory, based on stochastic analysis, have found a pragmatic way to deal with the “extensive form characteristics” of a continuous-time decision problem (see, e.g. [14, 13]). While extensive form theory is based on graph theory and refined partitions, stochastic analysis models exogenous information through filtrations on a given measurable space of scenarios. These are strictly less restrictive objects because $\sigma$-algebras need not be generated by partitions of the sample space. This works as if a one-shot lottery draw selects a scenario $\omega$ at the beginning, without communication, and then, as time progresses, more and more properties of $\omega$ become known to the decision makers. For instance, this could be knowledge about the realised path of a Brownian motion. From a decision-theoretic perspective, there is no disadvantage in using this weaker structure for modelling exogenous information.
@@
However, when it comes to the actual decision makers, the stochastic analysis-inspired approach is not compatible with the extensive form paradigm in the strict sense. This is because it does not explain strategies, randomisation, outcomes, or (dynamic, e.g. subgame-perfect) equilibria in terms of a decision tree-like object and choices locally available at moves. In this approach, “strategies” are typically defined as stochastic processes satisfying a certain measurability condition (e.g. progressive measurability) and such that a given stochastic differential equation, depending on them in a non-anticipating way, has a unique solution in some sense. This contradicts the problem’s own extensive form characteristics in the strict sense, due to the ex post restriction of the strategy space, which implies that the availability of a choice depends on future decisions (for a detailed discussion of this issue, see the introduction in [1]). The “outcome” is taken to be the solution to the differential equation, and “equilibria” (or “optimal controls”) are defined with implicit reference to the dynamic programming principle, but without an extensive form foundation. This not only fundamentally undermines the use of the equilibrium concept, but also introduces potential confusion, particularly regarding the decision-theoretic meaning of these “equilibria” (for example, concerning the notions of “closed” and “open loop” “equilibria” in stochastic differential games, see the discussion in [13, p. 72–76]), as well as the precise definition of subgames (for example, in stochastic timing games, see [41] for further discussion).
@@
Thus, there is room for an extensive form theory that models exogenous information through filtration-like objects, while modelling endogenous decision making with decision tree-like objects and partition-refining choices. This approach synthesises both frameworks in a productive way, which the author refers to as stochastic extensive forms. This is the aim of this project, with the present paper forming its first part. The theory can be used to construct extensive form games involving general stochastic processes as noise. Moreover, it provides a way to precisely characterise how stochastic control and differential games can be approximated by stochastic extensive form decision problems. For instance, it applies to the stochastic timing game, which is typically not at all formulated in extensive form (see, in particular, [41], where strategic form games are stacked according to a notion of “consistency”, which is justified a priori by an analogy to discrete time and a posteriori in [50] by a “discrete time with an infinitesimal grid” approximation argument reminiscent of [48, 49]). While a central motivation for this project stems from continuous-time problems, the theory is not confined to this domain. It is fundamentally about pushing the boundaries of extensive form theory as far as logically possible, with a focus set on probability. This makes it possible to formally establish links between uses of concepts such as equilibrium, which otherwise might seem ad hoc. It also allows for a clear outline of the decision-theoretic structure that arises naturally from the combination of these two ideas, distinguishing it from the additional structure sometimes added in modelling based on the modeller’s discretion or specific perspectives on a given problem. For example, stochastic extensive form theory is independent of the representation of noise through a “nature” agent’s virtual decision making, and in that sense, it is naturally stochastic.
@@
The first step in this endeavour, which is taken in the present paper, is to develop a theory of stochastic decision forests. These can be understood as forests of decision trees, where each tree corresponds to exactly one exogenous scenario, equipped with a similarity structure across trees, which identifies moves. Decision trees, as graph-theoretical objects, are the traditional base model for extensive form games and decision problems. However, as pointed out in [3], the refined partitions-based representation, using the set of all maximal chains, exhibits not only strong duality properties but also makes dynamic decision making amenable to the traditional decision-theoretic paradigm of choice under uncertainty: acts which assign a consequence to any state, following Savage’s framework ([45]). In this refined partitions approach, acts translate into strategies, mapping each move to a choice determined by a set of outcomes still possible at that move from the agent’s perspective. This is how uncertainty about choices (whether past or future choices of oneself or others, and present choices of opponents) – that is, endogenous information – is handled. The modelling of uncertainty about exogenous noise, or exogenous information, in this paper, however, is fundamentally different. It is the aforementioned similarity structure, consisting in so-called random moves, that can be equipped with a filtration-like object which dynamically reveals information about the realised exogenous scenario. This allows for general stochastic processes to model noise without running into outcome generation problems along that dimension. The adaptedness of strategies with respect to exogenous information can be based on a concept of adapted choices that is introduced in the final section.
@@
The fundamental motivation for this theory arises in a large class of concrete applications. Thus, its presentation is tightly accompanied by examples. While some of them remain rather pedagogical, a general class of stochastic decision forests based on time-indexed action paths is also constructed. A small set of readily verifiable axioms allows for many different time regimes, many different specifications of the outcomes (e.g. paths of timing games, also in the case of the scenario-dependent expiration of certain options), and general stochastic noise. Of course, it also includes the deterministic case. To the best of the author’s knowledge, this theoretical unification of action path-based decision problems with extensive form characteristics within a single framework is a second new contribution of this paper in its own right.
@@
Before concluding the introduction, three remarks remain to be made. First, this is the first of a series of three papers; thus, similar to the opening movement of a piano sonata, its first part is self-contained and addresses an independent research problem in its own right, but clearly prepares for and is closely related to the other two. Second, as pointed out by Aumann (see [6]), game theory arguably is “interactive decision theory”, while decision theory is, in a trivial sense, single-player game theory. Still, game theory and decision theory are not the same; the former is more concerned with the “interactive” aspect, while the latter focuses more on the “decision” facet. This paper is more concerned with the latter, so it primarily employs the corresponding terminology. However, when the context allows, “game” is used instead of “decision problem”, or vice versa, and similarly, the terms “decision maker”, “agent”, and “player” are used interchangeably. Note, however, that the concept of a decision maker is not formalised in this first paper; its use is purely motivational and of no structural importance at this stage. Finally, the relevant proofs of all new theorems, propositions, lemmata, and claims in examples can be found in the corresponding subsection of the appendix.
@@
  • $|M|$ = if $M$ is not an element of some real coordinate space $\mathbb{R}^{d}$, $d\in\mathbb{N}$, then this is the cardinality of the set $M$;
@@
  • $\mathscr{E}|_{D}=\{E\cap D\mid E\in\mathscr{E}\}$, for any $\sigma$-algebra $\mathscr{E}$ and any $D\in\mathscr{E}$;
  • $\mathscr{E}_{1}\vee\mathscr{E}_{2}$ = the smallest $\sigma$-algebra on $\Omega$ containing both $\mathscr{E}_{1}$ and $\mathscr{E}_{2}$, given a set $\Omega$ and sets $\mathscr{E}_{1},\mathscr{E}_{2}\subseteq\mathcal{P}(\Omega)$.
@@ 1 Decision forests
The basic object of classical extensive form decision and game theory is the decision tree. In the stochastic generalisation presented here, nature does not act as an agent taking decisions dynamically, but is simply replaced with a device “randomly” selecting the decision tree the “personal” agents follow during their decision making process. This implies that we consider decision forests, rather than decision trees. Although this term is not new (as testified by [43]), we consider it in the context of abstract decision theory, and more precisely within the – decision-theoretically natural – refined partitions framework, and in the aim of making this framework amenable to exogenous noise in the sense of general probability theory. This framework has originally been developed for trees, rather than forests, pioneered in [52, Section 8], and developed in much more generality in [3, 4] and subsequent papers, much of which is covered in the monograph [5]. In this section, we fix some order-theoretic language, present a definition and interpretation of decision forests within the refined partitions approach, and show that decision forests are nothing but forests of decision trees.
@@ 1.1 Order- and graph-theoretic conventions
The elements of $\mathcal{F}$ are called connected components of $(F,\geq)$. The maximum of a rooted tree $(T,\geq)$ is called the root. The roots of a forest $(F,\geq)$ are the roots of its connected components. A decision forest (decision tree) is a rooted forest (tree, respectively) $(F,\geq)$ such that all $x,y\in F$ with $x\neq y$ can be separated by some maximal chain $c\subseteq F$, that is, $c\cap\{x,y\}$ is a singleton. A move in a decision forest $(F,\geq)$ is a non-terminal node $x\in F$. In the present text, let us call a decision forest $(F,\geq)$ (everywhere) non-trivial iff some (any, respectively) root is a move.
@@ 1.2 Decision forests
The following definition is, formally, a transcription of the characterisation of (a subclass of) “game trees” in [3, Theorem 3]. So, although the object is formally not new, we look at it from a slightly different perspective which is why it is recalled here. First, it emphasises the fact that there can be multiple connected components, but restricts the attention to those [3]-“game trees” whose connected components are rooted. Second, the terminology in this text is different also in that it insists on the purely decision-theoretic aspect. What is called “decision forest (or tree) over a set” and “decision path” here, respectively, is called “game tree“ and “play” in [3]. Third, while [3]-“game trees” are defined via certain set-theoretic properties which can be rather easily verified in applications, and are then characterised via the so-called “representation by plays”, the present text perceives the decision-theoretic essence of decision forests over sets rather as being the duality between outcomes and nodes expressed by that representation, and thus uses it as the definition.
@@
Let $V$ be a set. A decision forest on $V$ is a $V$-poset $F$ such that:
@@
$F$ is called decision tree on $V$ iff, in addition, for all $x,y\in F$ there is $z\in F$ with $z\supseteq x\cup y$.
@@
Following [3], $V$ can be seen as the set of possible outcomes of the game, and the forest $(F,\supseteq)$ specifies how, as the decision problem is tackled dynamically, the set of realisable outcomes becomes smaller and smaller. Thus, elements of $F$ can be interpreted as nodes. See [3] for handy criteria on the set $F$ characterising the situation of it being a decision forest on $V$: essentially, this corresponds to $F$ being a “game tree” on $V$ in the sense of [3, Definition 4] such that all connected components of $(F,\supseteq)$ have maximal elements.
Condition 1.3.1 clarifies that the predecessors of any node $x\in F$ can be totally ordered and contain a root, and therefore constitute an account of the past with a beginning, a history. The set $W$ of maximal chains can be interpreted as the set of decision paths. Condition 1.3.2 says that possible outcomes and decision paths are in one-to-one correspondence, in such a way that a node is contained in some decision path iff the corresponding outcome is contained in that node. Even more is true: according to [3, Theorem 3, Corollary 2] and Remark 1.2, the bijection $f$ is uniquely determined. More precisely:
@@
Let $F$ be a decision forest on some set $V$ and $f\colon V\to W$ be a map as in Definition 1.3, where $W$ is the set of maximal chains in $(F,\supseteq)$. Then,
@@
Let $F$ be a decision forest on a set $V$. Then the poset $(F,\supseteq)$ is a decision forest.
@@
Then $F$ defines a decision forest on $V$, and $(F,\supseteq)$ is order-isomorphic to $(F_{0},\geq)$.
Rephrasing the “representation by plays” from [3], we call $F$ the representation by decision paths of $(F_{0},\geq)$. The definition of a decision forest on a set requires essentially that the operations from the two preceding proposition are, up to isomorphism, inverse to each other. See [3, Section 4] for more details on this.
Thus, there is a rigorous sense in that both of the mentioned descriptions are equivalent. While the graph-theoretical approach is graphically more convenient, the refined partitions approach is naturally aligned with Savage’s acts as a model for choice under uncertainty (see the introduction of [3] for a discussion). We note that, although phrased differently and with a different aim, all three preceding propositions and the duality concepts constitute one of the essential innovations of [3].